OpenStudy (mven):

Find the area of a circle circumscribed about an equilateral triangle whose side is 18 inches long.

OpenStudy (anonymous):

i will help you with this if you help me with mine pls

OpenStudy (anonymous):

If you draw this up and then draw from the center to the triangle angles you'll note that it splits the angles in two. We can solve for the radius using trig. We have a 30-120-30 triangle: sin(30)=r/18 18sin(30)=r=9 We can now solve for the circle are: Area=πr2=π(9)2=81π=

OpenStudy (anonymous):

That last part should say = 254.47

OpenStudy (mven):

thanks